$SPAD/src/input c05adf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
f:ASP1(F):=exp(-X)-X
a:SF:=0.0
b:SF:=1.0
eps:SF:=1.0e-5
eta:SF:=0.0
result:= c05adf(a,b,eps,eta,-1,f)

———-

References^{[1] nothing}

4